# Generalized Chern-Pontryagin models

### v1

@@ I Introduction
As it is known, various extensions of Einstein’s gravity have been proposed – those including modification in the pure gravity sector Sot ; DeFelice ; Noj and others involving extra fields, either scalar Brans ; Horndeski ; Koba , vector KosGra ; Jacob ; Del or even tensor ones (that is, so-called bimetric gravity). One of such extensions, that has attracted a strong interest in recent years, is the four-dimensional Chern-Simons modified gravity (CSMG) originally proposed in JaPi . In essence, the CSMG is the theory whose Lagrangian includes, besides of the scalar curvature, the Chern-Pontryagin topological term ${}^{*}RR\equiv\frac{1}{2}\epsilon^{\mu\nu\lambda\rho}R_{\mu\nu\alpha\beta}R_{\lambda\rho}^{\phantom{\lambda\rho}\alpha\beta}$, multiplied by an extra scalar $\vartheta$ called the Chern-Simons (CS) coefficient, i.e., the $\vartheta{}^{*}RR$ term (also called the CS term) is added. This term, being the natural four-dimensional extension of the 3-dimensional gravitational Chern-Simons term DJT , displays numerous interesting properties. First, it can break parity explicitly, by depending on the form of $\vartheta$; if $\vartheta$ is even under parity transformation, once ${}^{*}RR$ is odd under parity transformation. Also, it can break Lorentz symmetry at the action level, since one can rewrite the CS term as an interaction term between a background vector $v_{\mu}\equiv\partial_{\mu}\vartheta$ and the topological Chern-Simons current $K^{\mu}$ (for a discussion of other Lorentz/parity-breaking gravity models see KosGra ; Li ; Li2 ; Li3 ; Li4 ; Rao ). Second, it was checked that CSMG shares some solutions with the general relativity (GR), for example, the Schwarzschild metric. Conversely, one can cite, as a counterexample, the Kerr metric which is not a solution of CSMG, as discussed in Yunes ; Konno . In the further generalization of the CSMG, the so-called dynamical CSMG (DCSMG), the kinetic term for $\vartheta$ is introduced. A detailed discussion of consistency of various known metrics within the DCSMG is presented in Grumiller .
@@
In this paper, driven by $f(R)$ theories, we propose a new class of theories called generalized Chern-Pontryagin models, whose action, instead of being linear in ${}^{*}RR$, displays a generic functional dependency on this object and also the Ricci scalar. Certainly, this will allow for a richer framework to explore, among other issues, new parity-violating solutions of the modified field equations in the model. While many well-known metrics, including Schwarzschild and Gödel ones, actually, all metrics fulfilling the Pontryagin constraint ${}^{*}RR=0$, can be easily shown to be simultaneous solutions of GR and generalized Chern-Pontryagin theories. Then, it is interesting to examine the influence of the generalized Chern-Pontryagin term under small perturbations of these metrics within the modified field equations. Nonetheless, our main goal is to provide a natural mechanism to generate the four-dimensional CS term, which was originally proposed in analogy to the electromagnetism JaPi .
@@ II Generalized Chern-Pontryagin gravity models
It is worth stressing out that by taking the divergence of Eq.(4), we obtain the following constraint equation,
@@
where we have used the Bianchi identity, $\nabla^{\mu}G_{\mu\nu}=0$, and the conservation of the stress-energy tensor, $\nabla^{\mu}T^{(m)}_{\mu\nu}=0$, to find the aforementioned equation. Note that the divergence of the term in brackets in the r.h.s. of Eq. (6) resembles the Pontryagin constraint in the Chern-Simons modified theory of gravity (see f.e. Grumiller ). Another striking relation is obtained tracing Eq. (4). By doing so, we find
@@ V Summary
Acknowledgments. This work was supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq). PJP would like to thank the Brazilian agency CNPq for financial support (PQ–2 grant, process No. 307628/2022-1). The work by A. Yu. P. has been supported by the CNPq project No. 303777/2023-0.

### v2

@@ I Introduction
As it is known, various extensions of Einstein’s gravity have been proposed – those including modification in the pure gravity sector Sot ; DeFelice ; Noj ; Nojiri:2017ncd and others involving extra fields, either scalar Brans ; Horndeski ; Koba , vector KosGra ; Jacob ; Del or even tensor ones (that is, so-called bimetric gravity). One of such extensions, that has attracted a strong interest in recent years, is the four-dimensional Chern-Simons modified gravity (CSMG) originally proposed in JaPi . In essence, the CSMG is the theory whose Lagrangian includes, besides of the scalar curvature, the Chern-Pontryagin topological term ${}^{*}RR\equiv\frac{1}{2}\epsilon^{\mu\nu\lambda\rho}R_{\mu\nu\alpha\beta}R_{\lambda\rho}^{\phantom{\lambda\rho}\alpha\beta}$, multiplied by an extra scalar $\vartheta$ called the Chern-Simons (CS) coefficient, i.e., the $\vartheta{}^{*}RR$ term (also called the CS term) is added. This term, being the natural four-dimensional extension of the 3-dimensional gravitational Chern-Simons term DJT , displays numerous interesting properties. First, it can break parity explicitly, by depending on the form of $\vartheta$; if $\vartheta$ is even under parity transformation, once ${}^{*}RR$ is odd under parity transformation. Also, it can break Lorentz symmetry at the action level, since one can rewrite the CS term as an interaction term between a background vector $v_{\mu}\equiv\partial_{\mu}\vartheta$ and the topological Chern-Simons current $K^{\mu}$ (for a discussion of other Lorentz/parity-breaking gravity models see KosGra ; Li ; Li2 ; Li3 ; Li4 ; Rao ). Second, it was checked that CSMG shares some solutions with the general relativity (GR), for example, the Schwarzschild metric. Conversely, one can mention, as a counterexample, the Kerr metric which is not a solution of CSMG, as discussed in Yunes ; Konno ; Alexander:2009tp . In the further generalization of the CSMG, the so-called dynamical CSMG (DCSMG), the kinetic term for $\vartheta$ is introduced. A detailed discussion of the consistency of various known metrics within the DCSMG is presented in Grumiller , and various issues related to gravitational waves in metric DCSMG were studied in Nojiri:2019nar ; Nojiri:2020pqr . A metric-affine extension of the DCSMG with projective invariance has been developed Boudet , and propagation of gravitational waves within this theory has also been investigated Boudet2 ; Bombacigno . The Chern-Pontryagin topological invariant also appears naturally in the context of string/M-theory, specifically, compactifying $D=11$ M-theory down to $D=4$ effective gravity theories. In particular, one can cite the Starobinsky-Bel-Robinson (SBR) model Ketov:2022lhx ; Ketov:2022zhp , whose action involves quadratic Chern-Pontryagin terms, $({}^{*}RR)^{2}$, its cosmological aspects were studied in Ketov:2022lhx ; Ketov:2022zhp . Another interesting string-inspired model closely related to SBR model is the Einstein-Grisaru-Zanon effective gravity theory which has been investigated in CamposDelgado:2024jst .
@@
In this paper, driven by $f(R)$ theories, we propose a more generic class of theories called generalized Chern-Pontryagin models, whose action, instead of being linear or quadratic in ${}^{*}RR$, displays a generic functional dependency on this object and also the Ricci scalar. Certainly, this will allow for a richer framework to explore, among other issues, new parity-violating solutions of the modified field equations in the model. At the same time, many well-known metrics, including Schwarzschild and Gödel ones, actually, all metrics fulfilling the Pontryagin constraint ${}^{*}RR=0$, can be easily shown to be simultaneous solutions of GR and generalized Chern-Pontryagin theories. Then, it is interesting to examine the influence of the generalized Chern-Pontryagin term under small perturbations of these metrics within the modified field equations. Nonetheless, our main goal is to provide a natural mechanism to generate the four-dimensional CS term, which was originally proposed in analogy to the electromagnetism JaPi .
@@ II Generalized Chern-Pontryagin gravity models
Note however that the field equations involve higher-order derivative terms whose presence leads potentially to the emergence of ghost-like instabilities. At the perturbative level, it spoils the unitarity of the quantum field theory Hawking ; Antoniadis . At the classical one, it entails the arising of Ostrogradsky instabilities Ostrogradsky . A careful investigation of the number of the degrees of freedom by using a canonical analysis has been carried out for some higher-curvature gravity theories Crisostomi:2017ugk , including $f(\,^{*}RR)$ ones, where it was found that these theories possess at least one ghost-like mode. However, these difficulties are circumvented by following the methodology employed in effective field theories (EFT) Georgi , in which the ghost-like modes can be disregarded since they are suppressed by a typical high energy scale, rendering them heavy modes to propagate at the low energy limit. From now on, we shall treat the generalized Chern-Pontryagin models as EFTs.
It is worth stressing out that by taking the divergence of Eq.(2) [^1] For the pure $f(R)$ case this calculation was done in DeFelice ., we obtain the following constraint equation,
@@
where we have used the Bianchi identity, $\nabla^{\mu}G_{\mu\nu}=0$, and the conservation of the stress-energy tensor, $\nabla^{\mu}T^{(m)}_{\mu\nu}=0$, to find the aforementioned equation. Note that the divergence of the term in brackets in the r.h.s. of Eq.(6) resembles the Pontryagin constraint in the Chern-Simons modified theory of gravity (see f.e. Grumiller ). Another striking relation is obtained by tracing Eq. (4). By doing so, we find
@@ V Summary
Acknowledgments. This work was supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Paraiba State Research Foundation (FAPESQ-PB). PJP would like to thank the Brazilian agency CNPq for financial support (PQ–2 grant, process No. 307628/2022-1). The work by AYP has been supported by the CNPq project No. 303777/2023-0. PJP, AYP and JRN thank the Brazilian agency FAPESQ-PB for financial support (process No. 150891/2023-7).